OpenStudy (anonymous):

(2/7)r + 2r = (1/2)r + (17/2) Solve for r.

OpenStudy (anonymous):

I'm not really sure where to start :(

OpenStudy (anonymous):

ok, ill help,

OpenStudy (anonymous):

future warning, i might suddenly leave, but thats because my cp froze. sorry!

OpenStudy (oaktree):

I'll help then.

OpenStudy (anonymous):

That's okay :) I apprciate the help

OpenStudy (anonymous):

ok, oaktree, go right ahead.

OpenStudy (oaktree):

So, first you want to multiply each term by 2 to get rid of the fractions.

OpenStudy (oaktree):

You get 4/7r+4r=r+17

OpenStudy (oaktree):

Then multiply each term by 7 to get rid of the other fraction

OpenStudy (anonymous):

Oh! I see it's about getting rid of the fractions?

OpenStudy (oaktree):

And you get 4r+28r=7r+119

OpenStudy (oaktree):

No, it just makes it easier to solve.

OpenStudy (anonymous):

not exactly, but its helpful to simplify the, your goal is to isolate the r

OpenStudy (anonymous):

to simplify them*

OpenStudy (oaktree):

So then you can simplify and get 25r=119

OpenStudy (anonymous):

Oh! That makes sense. My professor kept telling me to multiply everything by 14

OpenStudy (oaktree):

That also works because it is really doing both at once, but it's simpler to do them in two steps to prevent mistakes.

OpenStudy (oaktree):

From 25r=119, you can easily get r=119/25.

OpenStudy (anonymous):

That's where I got confused! Thank you both sooo much!

OpenStudy (oaktree):

Which I don't believe can be simplified.

OpenStudy (oaktree):

So that's your final answer.

OpenStudy (oaktree):

You're welcome!

OpenStudy (oaktree):

Anything else I missed, mathfreak?

OpenStudy (oaktree):

You O.K. there?

OpenStudy (anonymous):

oaktree, is that right? mustn't you multiply it to the other side as well( simple an important rule, you can do something to one side of the equation, as long as you do the same to the other side.)

OpenStudy (oaktree):

What do you mean?

OpenStudy (oaktree):

Multiply what?

OpenStudy (oaktree):

If you mean the 2 and the 7, I did multiply both sides. Check my equation, maybe I messed it up somewhere.

OpenStudy (anonymous):

ya, i noticed that. did you inform her about it?

OpenStudy (anonymous):

sorry, this site crashes my cp,

OpenStudy (anonymous):

Yeah I saw it.